A narrow coil of 10 turns and area 4×10^-2m^2 is placed in a uniform magnetic field of flux density of 10^-2T so that the flux links the turns normally .calculate the accurate induced EMF in the coil if it is removed completely from the field in 0.55seconds

1
Expert's answer
2022-06-13T08:23:52-0400

Given:

"N=10"

"A=4*10^{-2}\:\rm m^2"

"B=10^{-2}\:\rm T"

"\Delta t=0.55\:\rm s"



"{\cal E}=-\frac{\Delta \Phi}{\Delta t}=-\frac{NA\Delta B}{\Delta t}"

"{\cal E}=-\frac{10*4*10^{-2}*(0-10^{-2})}{0.55}=7.3*10^{-3}\:\rm V=7.3\: mV"